OpenStudy (anonymous):

If the shadow of a tree is 14 m long and the shadow of a person who is 1.8 m tall is 4 m long, how tall is the tree? Answer to the nearest tenth. x = ?

OpenStudy (anonymous):

assuming he's standing adjacent to the tree, then both their shadows are cast horizontal on the ground forming right angled triangles; therefore: \[\tan \theta = (1.8/4) = (x/14)\] \[1.8 * 14 = 4x\] \[x = (1.8*14)/4\] x=?

OpenStudy (anonymous):

that makes sense :) so the answer would be 6.3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

you welcome, thanks